OpenStudy (bossimbacon):

how many solutions does this system of equations have y=-2x + 3 y=-2x-1 @texaschic14

OpenStudy (bossimbacon):

@texaschic101

pooja195 (pooja195):

Solve it using substitution :) y=-2x + 3 <---e1 y=-2x-1 <---e2 Plug e1 into e2 solve for x -2x+3=-2x-1 solve for x

Directrix (directrix):

y=-2x + 3 y=-2x - 1 The slopes of the two lines are the same. The lines have different y-intercepts. What is true about lines with the same slope?